OpenStudy (anonymous):

Let f(x) = (x-2)/(8) and g(x) = 8x + 2. Find (fog)(x)

OpenStudy (anonymous):

every place there is an x in f(x) put 8x+2

OpenStudy (unklerhaukus):

\[f(x)=\frac{x-2}8\qquad\qquad g(x)=8x+2\] \[(f\circ g)x=f(g(x))=f(8x+2)=\frac{(8x+2)-2}8=\]

OpenStudy (anonymous):

make sure to simply

OpenStudy (anonymous):

is fog(x) like saying f of g of x,

OpenStudy (unklerhaukus):

exactly @amcdowns